# DyFFPAD: Dynamic Fusion of Convolutional and Handcrafted Features for Fingerprint Presentation Attack Detection

### v2

@@ 5.2 Comparative Analysis
In this section, findings of the proposed method are compared with state-of-the-art methods as per the spoofs fabricated with known and unknown materials. A detailed comparative analysis is provided in the following subsections.
TABLE V: Comparison with state-of-the-art methods on LivDet 2015
@@
Accuracy
Accuracy
Accuracy
(Digital Persona)
Accuracy
@@
Kim et al. [18] - - - - 86.39
Uliyan et al. [35] 95.00 - - - 95.00
Sharma et al. [32] 98.07 95.7 94.16 95.22 95.78
Jung et al. [16] 98.60 96.20 90.50 95.80 95.27
Sharma et al. [31] 95.0 96.7 97.2 95.50 96.10
Xia et al.[37] 89.18 95.47 86.28 90.36 90.32
Spinoulas et al. [33] 98.10 98.56 94.80 96.80 97.11
LivDet 2015 Winner [23] 98.10 95.40 93.72 94.36 95.39
Yuan et al.[41] 96.54 95.32 93.2 93.76 94.70
Zhang et al. [43] 97.01 97.81 95.42 97.02 96.82
@@
5.2.1 Comparison with state-of-the-art methods on LivDet 2015 database
The LivDet 2015 database includes spoof samples fabricated using both known and unknown spoofing materials. Table V provides a detailed comparison of the proposed method’s performance with state-of-the-art methods validated on this database. Table V clearly shows that the proposed method outperforms the methods discussed in [32], [17], [23], [35], [31], [37], [39], and [18]. while it is comparable to the methods discussed in [33] and [43]. The proposed method achieves an accuracy of 99.29% which is the best as compared with all other methods validated on crossmatch sensor. The proposed method outperforms most other methods, achieving an overall classification accuracy of 96.10
5.2.2 Comparison with state-of-the-art methods on LivDet 2017 database
The performance of the proposed method is also compared with methods validated on the LivDet 2017 database. Notably, the training and testing spoof samples within this database are fabricated using different spoofing materials. Our method successfully demonstrates robust classification capabilities in distinguishing between live and spoof samples, achieving commendable accuracy. Table VI shows that the proposed method outperforms the methods discussed in [5], [4], [43], and [12] when evaluated on the fingerprint samples captured with orcanthus and digital persona sensors while it is comparable with [28] on orcanthus sensor. The proposed method achieves an accuracy of 96.00% on greenbit sensor which is better than of [12], [43] and is comparable with [4]. The proposed method also outperforms the state-of-the-art methods in terms of overall classification accuracy of 96.49%. This comparison reveals that the dynamic ensemble of handcrafted and convolutional features is able to detect spoofs regardless of the material used for fabrication.
TABLE VI: Comparison with state-of-the-art methods on LivDet 2017 database
@@ 5.2.2 Comparison with state-of-the-art methods on LivDet 2017 database
Accuracy
Accuracy
(Digital Persona)
Accuracy
@@
Zhang et al. [43] 93.93 92.89 95.20 94.00
Rai et al. [28] 96.53 93.64 94.74 94.97
Chugh et al. [5] 95.01 95.20 97.42 95.88
Gonzalez et al. [12] 94.38 95.08 94.54 94.66
Chugh et al. [4] 94.51 95.12 96.68 95.43
@@
5.2.3 Comparison with state-of-the-art methods on LivDet 2019 database
Table VII compares the performance of the proposed model with state-of-the-art methods evaluated on the LivDet 2019 database. This comparison reveals that the proposed method achieves an accuracy of 91.57%, outperforming the approaches outlined in [4] as well as the FPAD algorithms, including JungCNN, JWL LivDet, and ZJUT DET, when tested on samples captured with digital persona sensor. For the orcanthus sensor, the proposed method outperforms the majority of the methods with a classification accuracy of 98.32%. The overall performance comparison indicates that the proposed method outperforms all other methods, achieving a classification accuracy of 95.99
The comparative analysis mentioned above concludes that the proposed method consistently outperforms state-of-the-art methods in the intra-sensor paradigm of FPAD, regardless of whether the spoof samples are fabricated with known or unknown materials. As compared to standard CNN-based techniques, the dynamic ensemble of DNN empowered with handcrafted features and CNN enables them to learn their parameters better.
@@ 5.2.3 Comparison with state-of-the-art methods on LivDet 2019 database
TABLE VII: Comparison with state-of-the-art methods on LivDet 2019 database
@@
Accuracy
Accuracy
(Digital Persona)
Accuracy
Average
Chugh et al. [4] 97.50 83.64 99.73 93.62
Jung CNN [25] 99.13 81.23 99.06 93.14
JWL LivDet [25] 97.45 88.86 99.20 95.17
ZJUT Det A [25] 97.50 88.77 99.20 95.16
@@ 5.3 Evaluation of DyFFPAD in High-Security Systems
In the context of high-security systems, it is imperative to evaluate the performance of an FPAD model comprehensively. While the model’s fundamental objective is indeed to minimize the APCER, BPCER, and ACE, its assessment should extend beyond these metrics to ensure its efficacy in safeguarding high-security environments. In this study, we present the results of the proposed model using the Detection Error Trade-off (DET) curve. The DET curve is a graphical representation of the error rates achieved by a binary classification system as the classification threshold value is varied. The DET curves for all datasets from the LivDet 2015, 2017, and 2019 databases are shown in Fig. 6. As shown in Fig. 6 it is evident that the proposed model achieves a BPCER of less than 1% to achieve APCER of 1% when evaluated on crossmatch, and it is in the range of 15% - 40% for biometrika, greenbit and digital persona sensors of LivDet 2015 database. On the LivDet 2017 database, the model is able to restrict BPCER in the range of 7% - 18% when testing spoof samples are obtained using unknown spoof materials. Similarly, the model maintains a BPCER of less than 5% for greenbit and orcanthus sensors of LivDet 2019 database.
@@ 6 Conclusion
  • The publicly available databases used for the validation of this work can be accessed via the following link: [https://livdet.org/registration.php].

### v5

@@ 5.1 Comparative Analysis
In this section, state-of-the-art methods are compared with the results of the proposed method in different evaluation conditions. A detailed comparative study is provided in the following subsections.
Table 5: Comparison with state-of-the-art methods on LivDet 2015
@@
Crossmatch
Greenbit
Digital Persona
Biometrika
@@
Uliyan et al. (2020) 95.00 - - - 95.00
Sharma and Dey (2019) 98.07 95.7 94.16 95.22 95.78
Jung and Heo (2018) 98.60 96.20 90.50 95.80 95.27
Sharma and Selwal (2021) 95.0 96.7 97.2 95.50 96.10
Xia et al. (2020) 89.18 95.47 86.28 90.36 90.32
LivDet 2015 Winner Mura et al. (2015) 98.10 95.40 93.72 94.36 95.39
Yuan et al. (2020) 96.54 95.32 93.2 93.76 94.70
Zhang et al. (2019) 97.01 97.81 95.42 97.02 96.82
Rai et al. (2024) 97.51 97.63 93.50 96.18 96.20
Rubab Mehboob (2023) 97.9 97.48 88.96 97.9 95.5
Rai and Dey (2024) 95.01 96.95 93.32 96.00 95.32
Kaur (2024) - - - 96.49
@@
5.1.1 Comparative study on LivDet 2015 database
Table 5 provides a detailed comparison of the proposed method’s performance with state-of-the-art methods. The results presented in Table 5 clearly show that the proposed method achieves an accuracy of 99.29% which is the best as compared with all other methods validated on crossmatch sensor dataset while it performs better than all other methods except Rubab Mehboob (2023) and Zhang et al. (2019) while being validated on biometrika dataset. In terms of overall performance, The proposed method outperforms the methods presented in Sharma and Dey (2019), Jung et al. (2019), Mura et al. (2015), Uliyan et al. (2020), Sharma and Selwal (2021), Xia et al. (2020), Yuan et al. (2022), and Rai and Dey (2024) while it is comparable to the methods discussed in Kaur (2024), Rai et al. (2024), and Zhang et al. (2019).
5.1.2 Comparative study on LivDet 2017 database
The performance of the proposed method is also compared with methods validated on the LivDet 2017 database. Our method successfully demonstrates robust classification capabilities in distinguishing between live and spoof samples, achieving commendable accuracy. The comparison shown in Table 6 indicate that our method is able to perform better as compared with the methods discussed in Chugh and Jain (2021), Chugh et al. (2018), Zhang et al. (2019), Rai et al. (2025), and González-Soler et al. (2021) for orcanthus and digital persona sensors while it is comparable with Rai et al. (2024) on orcanthus sensor. The proposed method achieves an accuracy of 96.00% on greenbit dataset which is better than of González-Soler et al. (2021), Zhang et al. (2019), Rai et al. (2025) and is comparable with Chugh et al. (2018). The overall classification accuracy achieved by the proposed method is 96.49% which is better than all other methods. This comparison reveals that the dynamic ensemble of handcrafted and convolutional features is able to detect spoofs irrespective of the material used for their fabrication.
Table 6: Comparison with state-of-the-art methods on LivDet 2017 database
@@ 5.1.2 Comparative study on LivDet 2017 database
Orcanthus
Digital Persona
Greenbit
@@
Zhang et al. (2019) 93.93 92.89 95.20 94.00
Chugh and Jain (2021) 95.01 95.20 97.42 95.88
González-Soler et al. (2021) 94.38 95.08 94.54 94.66
Chugh et al. (2018) 94.51 95.12 96.68 95.43
Rai et al. (2024) 96.53 93.64 94.74 94.97
Rai et al. (2025) 95.94 95.40 93.79 95.05
@@
5.1.3 Comparative study on LivDet 2019 database
Table 7 compares the performance of the proposed model on the LivDet 2019 database. This comparison reveals that the proposed method achieves an accuracy of 91.57%, outperforming the approaches outlined in Chugh et al. (2018) as well as the FPAD algorithms, including JWL LivDet, JungCNN, and ZJUT DET, when tested on samples captured with digital persona sensor. For the orcanthus sensor, the proposed method outperforms all other methods with a classification accuracy of 98.32% while it remains second after González-Soler et al. (2021) with an accuracy of 91.57% on digital persona dataset. The overall performance comparison indicates that the proposed method outperforms all other methods except, González-Soler et al. (2021), achieving a classification accuracy of 95.99
The comparative analysis mentioned above concludes that the proposed method consistently outperforms state-of-the-art methods in the intra-sensor paradigm of FPAD, for both known or unknown materials scenario. As compared to standard CNN-based techniques, the dynamic ensemble of DNN empowered with handcrafted features and CNN enables them to learn their parameters better.
@@ 5.1.3 Comparative study on LivDet 2019 database
Table 7: Comparison with state-of-the-art methods on LivDet 2019 database
@@
Orcanthus
Digital Persona
Greenbit
Average (%)
Chugh et al. (2018) 97.50 83.64 99.73 93.62
JWL LivDet Orrù et al. (2019) 97.45 88.86 99.20 95.17
ZJUT Det A Orrù et al. (2019) 97.50 88.77 99.20 95.16
Rai et al. (2024) 97.55 81.55 98.71 92.9
Rai et al. (2025) 97.72 90.59 97.31 95.20
González-Soler et al. (2021) 97.21 97.68 93.63 96.17
@@ 5.2 Evaluation of DyFFPAD in High-Security Systems
In the context of high-security systems, it is imperative to evaluate the performance of an FPAD model comprehensively. While the model’s fundamental objective is indeed to minimize the APCER, BPCER, and ACE, its assessment should extend beyond these metrics to ensure its efficacy in safeguarding high-security environments. In this research, we report the findings of the suggested model using the Detection Error Trade-off (DET) curve, which is a graphical depiction of the error rates produced by a binary classification system when the classification threshold value is varied. The DET curves for all datasets from the LivDet 2015, 2017, and 2019 databases are shown in Fig. 6. As shown in Fig. 6 it is evident that the proposed model achieves a BPCER of less than 1% to achieve APCER of 1% when evaluated on crossmatch, and it is in the range of 15% - 40% for biometrika, greenbit and digital persona sensors of LivDet 2015 database. On the LivDet 2017 database, the model is able to restrict BPCER in the range of 7% - 18% when testing spoof samples are obtained using unknown spoof materials. Similarly, the model maintains a BPCER of less than 5% for greenbit and orcanthus sensors of LivDet 2019 database.
@@ 6 Conclusion
  • The publicly accessible datasets used to validate this work can be obtained at [https://livdet.org/registration.php].